# Towards Principled Superhuman AI for Multiplayer Symmetric Games

### v1

@@ 6 Experiments
In this section, we focus on the scenario where one AI agent competes against $n-1$ human players who play the identical meta-strategy. We aim to answer: (Q1) Can existing algorithmic frameworks in previous superhuman AI systems consistently secure an equal share when competing against human players in games? If not, what are the patterns of the failure cases? (Q2) Are these trained agents exploitable by adversarial opponents? We design the following two games to compare our proposed methods and other baselines based on self-play.
@@
We first consider the standard majority vote game (Example 1) with $3$ players. A payoff of 0.5 is evenly distributed among the majority, and a payoff of -0.5 is evenly distributed among the minority. Note that everyone receives 0 if choosing the same action. Here the mixed strategy of the human population is assumed to be $\pi_{{\text{human}}}=[0.49,0.51]$. It can be seen that [1, 0], [0, 1], and [1/2, 1/2] are all NE.
@@
In each round, players simultaneously choose an action from actions $A$, $B$, or $C$. Let $N$ be the total number of players and $n_{A}$ be the number of agents choosing action $A$, We define the game rule as:
@@
where the rule $i\succ j\succ k$ intuitively means that action $i$ dominates both $j$ and $k$, and action $j$ dominates $k$. SDG is designed so that $C$ is a dominated action when there is a reasonable number of players taking action $A$, but a dominating action otherwise. Concretely, for $i\succ j\succ k$, we use the payoff defined as:
$r_{i}$ $=\mathbb{I}[n_{j}+n_{k}>0],$
$r_{j}$ $=\mathbb{I}[n_{k}>0]-\mathbb{I}[n_{j}+n_{k}>0]\frac{n_{i}}{n_{j}+n_{k}},$
$r_{k}$ $=-\mathbb{I}[n_{j}+n_{k}>0]\frac{n_{i}}{n_{j}+n_{k}}-\mathbb{I}[n_{k}>0]\frac{n_{j}}{n_{k}}.$
This payoff design guarantees that SDG is a zero-sum game. Throughout our experiments, we choose $N=30$ and the human meta-strategy $\pi_{{\text{human}}}=[0.399,0.6,0.001]$ (in the order of $A$, $B$, and $C$). Note that while this game has an NE policy [0, 0, 1], its utility is negative against all other players playing $\pi_{{\text{human}}}$.
@@ 6.1 Learning algorithms
To better focus on the game-theoretical part of the algorithms, we idealize the process of imitation learning by assuming that the agent has access to the human population policy $\pi_{{\text{human}}}$. We aim to compare our solution concepts — best response to behavior cloning (BR_BC), with three meta-algorithms adopted previously to build superhuman AI agents: self-play from scratch (SP_scratch) Brown and Sandholm, (2019), self-play from behavior cloning (SP_BC) Li et al., (2020), and self-play from behavior cloning with regularization towards human behavior (SP_BC_reg) Jacob et al., (2022). While these AI systems further implement multi-step lookahead with a few additional techniques, many of them only apply to sequential games, rather than basic normal-form games. Here, we focus on the comparison of the high-level meta-algorithms.
@@
We use the Hedge algorithm as the backbone for self-play. Compared with the vanilla SP_scratch which is initialized with the uniform action distribution, SP_BC starts from the human policy $\pi_{{\text{human}}}$, while SP_BC_reg (Jacob et al.,, 2022) further regularizes the KL divergence between the learned policy and $\pi_{{\text{human}}}$ during training. The Hedge algorithm also provides a natural way to learn the best response in BR_BC. We choose the learning rate for the Hedge algorithm based on theoretically optimal value, and choose the regularization parameter according to Jacob et al., (2022). We refer readers to Appendix E for more details.
@@ 6.2 Results
To answer Q1 and Q2, we evaluate the utility of the learned strategy against the human population, i.e., $U_{1}(\pi_{{\text{agent}}},\pi_{{\text{human}}}^{\otimes(n-1)})$, as well as the exploitability of $\pi_{{\text{agent}}}$, i.e., $\min_{\pi}U_{1}(\pi_{{\text{agent}}},\pi^{\otimes(n-1)})$. To measure the utility, we evaluate the payoff of the agent’s converged policy by Monte Carlo methods with $3\times 10^{5}$ games, and report the mean and standard deviation of 10 runs. As for the exploitability, we pick the best exploiter policy within 100 runs and report the corresponding utility.
@@
Before comparing the utility and exploitability, we first check the convergence policy for each algorithm:
(1) MV: We notice that BR_BC consistently converges to the policy [0,1], while self-play variants approach different equilibria across different runs, as summarized in Table 1. Therefore, all self-play variants can converge to a suboptimal strategy [1,0] with a non-negligible probability and suffer a negative payoff against the human population policy [0.49,0.51]. Based on this analysis, we report the utility and exploitability of the convergence policy [0,1] for BR_BC, and the suboptimal policy [1,0] for algorithms based on self-play.
(2) SDG: We notice that self-play algorithms consistently converge to the policy [0, 0, 1], while BR_BC converges to the policy [0, 1, 0]. Based on this analysis, we report the utility and exploitability of the convergence policy [0, 0, 1] for algorithms based on self-play, and the policy [0, 1, 0] for BR_BC.
MV SP_scratch SP_BC SP_BC_reg ($\lambda=10^{-5}/10^{-4}/10^{-3}/10^{-2}$)
$[1,0]$ 52 48 48
$[0,1]$ 48 52 52
Table 1: The distribution of convergence policy for self-play algorithms in 100 runs. For SP_BC_reg, we report results with different regularization coefficients $\lambda$.
MV SP_scratch / SP_BC / SP_BC_reg BR_BC
Utility (×$\times 10^{-2}$) -0.50 $\pm$ 0.05 0.52 $\pm$ 0.05
Exploitability -0.50 $\pm$ 0.00 -0.50 $\pm$ 0.00
SDG SP_scratch / SP_BC / SP_BC_reg BR_BC
@@
Table 2: The utility and exploitability of each algorithm. Particularly, for MV, as self-play algorithms converge to different solutions, we evaluate the suboptimal one with the negative utility, i.e., [1, 0]. The results show that in two well-designed games, all algorithms except for BR_BC achieve negative payoffs even against non-adaptive humans, and all the learned policies can be exploited.
@@
We summarize the results in Table 2, which show that even in these two simple cases, none of the self-play algorithms can consistently secure an equal share in symmetric constant-sum games. This undesirable behavior persists even without humans making any adaptations! Moreover, based on these two games, we further conclude two potential failure modes of self-play algorithms: (1) For games with multiple NE, self-play methods based on no-regret algorithms may converge to different NEs according to different initial policies. When human policy (the initial policy for SP_BC) lies close to the boundary of the convergence basins of two different NEs, self-play algorithms will have a non-zero probability to converge to both of them due to the statistical randomness in the game. It is likely one of the two NEs is highly suboptimal against $\pi_{\text{human}}$. Admittedly, this statistical randomness can be, to some extent, mitigated by picking a sufficiently small learning rate. However, the closer the initial policy to the boundary, the smaller the learning rate we need, which quickly makes the learning rate too small to converge within a reasonable number of samples in practice. (2) Even for a game with a single NE, a carefully designed game structure can result in this NE yielding a negative utility when compared to a specific human policy, and hence jailbreak all self-play variants. The aforementioned failure modes highlight a significant limitation in self-play variants’ capability to generalize effectively to diverse and complex game scenarios. In contrast, the solution BR_BC indicated by our theory consistently outperforms human policies and receives much higher utility. On the exploitability side, all learned policies can be exploited easily by adversarial opponents.
@@ 7 Conclusion
This paper takes the initial step towards addressing unique challenges in multiplayer games by investigating symmetric zero-sum normal-form games, which commonly appear in our daily lives. We clarify a number of important solution concepts for multiplayer games, and investigate the behaviors and limitations of many existing meta-algorithms that were deployed by previous state-of-the-art AI systems for games. We hope our results promote further research in principled methodologies and algorithms that lead to general superhuman AI for a wide range of multiplayer games.
@@ Acknowledgement
The authors would like to thank Haifeng Xu for helpful discussions. This work is supported by Office of Naval Research N00014-22-1-2253.
Appendix A Proofs for Section 4
A.1 $3$-player majority and minority game
In this paper, we define the 3-player majority game as a symmetric zero-sum game with action space $\mathcal{A}:=\{0,1\}$ and the payoff function given by:
@@ A.1 $3$-player majority and minority game
A.2 Proof of Proposition 4.3
@@
First of all, we show that
$\max_{x_{1}}\min_{x_{2},\cdots,x_{n}}U_{1}(x_{1},\cdots,x_{n})\leq\min_{x_{2},\cdots,x_{n}}\max_{x_{1}}U_{1}(x_{1},\cdots,x_{n})\leq\min_{x}\max_{x_{1}}U_{1}(x_{1},x^{\otimes n-1})=0,$
where all the inequalities can be strict.
@@
The second inequality is straightforward due to a restriction on the minimization constraints. In the sequel, we prove the equation via contradiction. Note that by choosing $x_{1}=x$, we can show that
$\min_{x,\cdots,x}\max_{x_{1}}U_{1}(x_{1},x,\cdots,x)\geq 0.$
@@
which contradicts with the assumption that the game is zero-sum and symmetric. Consequently, we prove the equation.
In Claim 4.1, we have proved that the second inequality can be strict. To show the first inequality can be strict, we consider the $3$-player majority vote. Suppose $3$ players adopt the mixed strategies $(\alpha_{1},1-\alpha_{1})$, $(\alpha_{2},1-\alpha_{2})$ and $(\alpha_{3},1-\alpha_{3})$, respectively. It then holds that
$U_{1}(x_{1},x_{2},x_{3})$
$=U_{1}(\alpha_{1},\alpha_{2},\alpha_{3})$
$=\alpha_{1}\left(-2(1-\alpha_{2})(1-\alpha_{3})+\alpha_{2}(1-\alpha_{3})+\alpha_{3}(1-\alpha_{2})\right)$
$+(1-\alpha_{1})\left(-2\alpha_{2}\alpha_{3}+\alpha_{2}(1-\alpha_{3})+\alpha_{3}(1-\alpha_{2})\right).$
$\max_{\alpha_{1}}\min_{\alpha_{2},\alpha_{3}}U_{1}(\alpha_{1},\alpha_{2},\alpha_{3})\leq\max_{\alpha_{1}}\min\{-2\alpha_{1},-2(1-\alpha_{1})\}=-1.$
Note that
$=\min_{\alpha_{2},\alpha_{3}}\max\{-2(1-\alpha_{2})(1-\alpha_{3})+\alpha_{2}(1-\alpha_{3})+\alpha_{3}(1-\alpha_{2}),-2\alpha_{2}\alpha_{3}+\alpha_{2}(1-\alpha_{3})+\alpha_{3}(1-\alpha_{2})\}$
$=\min_{\alpha_{2},\alpha_{3}}\max\{3(\alpha_{2}+\alpha_{3})-4\alpha_{2}\alpha_{3}-2,\alpha_{2}+\alpha_{3}-4\alpha_{2}\alpha_{3}\}$
$=0.$
Thus, we show that $\max_{\alpha_{1}}\min_{\alpha_{2},\alpha_{3}}U_{1}(\alpha_{1},\alpha_{2},\alpha_{3})<\min_{\alpha_{2},\alpha_{3}}\max_{\alpha_{1}}U_{1}(\alpha_{1},\alpha_{2},\alpha_{3})$.
We then show that
$\max_{x_{1}\in\Delta(\mathcal{A})}\min_{x\in\Delta(\mathcal{A})}U_{1}(x_{1},x^{\otimes n-1})\leq\min_{x\in\Delta(\mathcal{A})}\max_{x_{1}\in\Delta(\mathcal{A})}U_{1}(x_{1},x^{\otimes n-1})=0,$
where the inequality can be strict.
@@
In Claim 4.2, we have proved that the inequality can be strict. ∎
A.3 Proof of Proposition 4.4
@@
Appendix B Proofs for Section 5.1
B.1 Proof of Theorem 5.1

### v2

@@ 6 Experiments
In this section, we focus on the scenario where one learning agent competes against $n-1$ opponents who play the identical meta-strategy. For simplicity, we restrict ourselves to the setting of fixed opponents. We aim to answer: (Q1) Can existing algorithmic frameworks in previous superhuman AI systems consistently secure an equal share under this favorable setting? If not, what are the failure cases? (Q2) Are these trained agents exploitable by adversarial opponents? We design the following two games to compare our algorithm with prior self-play-based algorithms.
@@
We first consider the standard 3-player majority vote game (Example 1). It is not hard to see that $[1,0]$, $[0,1]$, and $[1/2,1/2]$ are all NEs, where $[p,1-p]$ denotes the mixture strategy that takes the first action with probability $p$ and the second action with probability $1-p$. We fix the opponents’ meta-strategy $y_{\text{meta}}=[0.49,0.51]$ for all rounds.
@@
In each round, players simultaneously choose an action from the set $\{A,B,C\}$. Let $n$ be the total number of players and $n_{A}$ be the number of agents choosing action $A$, We define the game rule as:
@@
where the rule $i\succ j\succ k$ intuitively means that action $i$ dominates both $j$ and $k$, and action $j$ dominates $k$. SDG is designed so that $C$ is a dominated action when there is a reasonable number of players taking action $A$, but a dominating action otherwise. Concretely, for $i\succ j\succ k$, we assign the following payoff $(r_{i},r_{j},r_{k})$ to players taking actions $(i,j,k)$ respectively, where:
$\textstyle r_{i}$ $=\mathbb{I}[n_{j}+n_{k}>0],\quad r_{j}=\mathbb{I}[n_{k}>0]-\mathbb{I}[n_{j}+n_{k}>0]\cdot n_{i}/(n_{j}+n_{k}),$
$r_{k}$ $=-\mathbb{I}[n_{j}+n_{k}>0]\cdot n_{i}/(n_{j}+n_{k})-\mathbb{I}[n_{k}>0]\cdot n_{j}/n_{k}.$
This payoff design guarantees that SDG is a symmetric zero-sum game. Throughout our experiments, we choose $n=30$ and pick the fixed meta-strategy of the opponents $y_{\text{meta}}=[0.399,0.6,0.001]$ (in the order of action $A,B,C$) for all rounds. Note that while this game has an NE strategy [0, 0, 1], its utility is negative against our chosen meta-strategy $y_{\text{meta}}$.
@@ 6.1 Learning algorithms
To better focus on the key game-theoretic property of the algorithms, we idealize the process of imitation learning by assuming that the learning agent has already learned (i.e., has direct access to) the meta-strategy $y_{\text{meta}}$ by the opponents. In this setting, according to Theorem 5.1, our theoretical framework suggests to directly run the Hedge algorithm (Hedge) against opponents who play the meta-strategy. We compare our algorithm against three meta-algorithms adopted by prior state-of-the-art AI systems in practice: (1) self-play from scratch (SP_scratch) (Brown & Sandholm, 2019); (2) self-play initialized from behavior cloning (SP_BC) (Li et al., 2020), and (3) self-play initialized from behavior cloning with regularization towards the meta-strategy (SP_BC_reg) (Jacob et al., 2022). While these AI systems further implement multi-step lookahead with a few additional techniques, many of them only apply to sequential games, not the basic normal-form games. Here, we focus on the comparison of the high-level game-theoretic meta-algorithms.
@@
We also use the Hedge algorithm as the optimizer for the self-play algorithm to update its strategy. We choose the learning rate for the Hedge algorithm based on theoretically optimal value and choose the regularization parameter according to Jacob et al. (2022). We refer readers to Appendix D for more details.
@@ 6.2 Results
To answer Q1 and Q2, we evaluate the utility of the learned strategy $\hat{x}$ against the pre-specified meta-strategy $y_{\text{meta}}$, i.e., $U_{1}(\hat{x},y_{\text{meta}}^{\otimes(n-1)})$, as well as the exploitability of the learner $\hat{x}$, i.e., $\min_{y}U_{1}(\hat{x},y^{\otimes(n-1)})$. To measure the utility, we evaluate the payoff of the agent’s converged strategy by Monte Carlo methods with $3\times 10^{5}$ games and report the mean and standard deviation of 10 runs. As for the exploitability, pick the best exploiter strategy within 100 runs, and report the payoff of the learner against that exploiter.
@@
We first check the convergence for each algorithm in both games:
MV: We report the limiting solution each algorithm converges to within 100 runs, as summarized in Table 1. Our algorithm (Hedge) consistently converges to the good strategy $[0,1]$. All self-play variants have significant probability converge to the bad strategy $[1,0]$, which has a negative utility against the chosen meta-strategy $y_{\text{meta}}=[0.49,0.51]$.
SDG: We report that while our Hedge algorithm converges to the strategy $[0,1,0]$, all self-play variants consistently converge to the strategy $[0,0,1]$.
Strategy SP_scratch SP_BC SP_BC_$10^{-5}$ SP_BC_$10^{-4}$ SP_BC_$10^{-3}$ SP_BC_$10^{-2}$ Hedge
$[1,0]$ $52\%$ $48\%$ $50\%$ $46\%$ $42\%$ $42\%$ $0\%$
$[0,1]$ $48\%$ $52\%$ $50\%$ $54\%$ $58\%$ $58\%$ $100\%$
Table 1: The distribution of mixed strategies to which different self-play algorithms converge in MV. We use SP_BC_$\lambda$ as the short name of SP_BC_reg with regularization coefficients $\lambda$.
MV SP_scratch / SP_BC / SP_BC_reg Hedge
Utility (×$\times 10^{-2}$) -1.00 $\pm$ 0.09 1.03 $\pm$ 0.10
Exploitability -1.00 $\pm$ 0.00 -1.00 $\pm$ 0.00
SDG SP_scratch / SP_BC / SP_BC_reg Hedge
@@
Table 2: The utility and exploitability of each algorithm. Particularly, for MV, as self-play algorithms converge to two different solutions with roughly equal probability, we evaluate the utility of the worse converged solution of the two to reflect the performance in the worst case.
@@
We summarize the results in Table 2, which show that even in these two simple symmetric zero-sum games, none of the self-play algorithms can consistently secure a non-negative payoff, i.e., an equal share, in the worst case. This undesirable behavior persists even without opponents making any adaptations! Moreover, based on these two games, we further conclude two potential failure modes of self-play algorithms: (1) For games with multiple NEs, such as MV, self-play methods may converge to different NEs based on different initialization. When the opponents’ meta-strategy (i.e., the initial strategy for SP_BC) lies close to the boundary of the convergence basins of two different NEs, self-play algorithms will have a non-zero probability of converging to both of them due to the statistical randomness in the game. It is likely one of the two NEs is undesirable against the meta-strategy $y_{\text{meta}}$. (2) For games with a single NE, self-play algorithms are still very likely to be attracted to this equilibrium. A carefully designed game structure can result in this NE yielding a negative utility against the chosen meta-strategy, and hence jailbreak all self-play variants. The aforementioned failure modes highlight a significant limitation of self-play variants when being applied to diverse and complex multiplayer games. In contrast, the principled algorithm according to our theory consistently beats the meta-strategy of the opponents, receives a much higher payoff, and secures an equal share. Regarding exploitability, all learned strategies can be easily exploited by adversarial opponents.
@@ 7 Conclusion
Unlike in the two-player zero-sum games, standard equilibria are no longer always the suitable solution concept for multiplayer games. They are non-unique and lack meaningful guarantees in general scenarios. This paper establishes a new theoretical framework that provides the solution concepts and the principled algorithms for multiplayer games from the unique angle of achieving equal share. We hope our results serve as the first step toward further research on principled methodologies and algorithms for multiplayer games.
@@ Acknowledgement
The authors would like to thank Haifeng Xu for the helpful discussions. This work is supported by Office of Naval Research N00014-22-1-2253.
Appendix A Extended Preliminaries
A.1 No-regret learning
No-regret learning is a commonly adopted strategy in game theory to find equilibrium solutions. We consider a $T$-step learning procedure, where for each round $t\in[T]$: (1) the agent picks a mixed strategy $\mu^{t}$ over $\mathcal{A}$, (2) the environment picks an adversarial loss $\ell_{t}\in[0,1]^{|\mathcal{A}|}$. The expected utility for $t$-th round is defined as $-\langle\mu^{t},\ell_{t}\rangle$. To measure the performance of a particular algorithm, a common approach is to consider regret, where the algorithm’s performance is compared against the single best action in hindsight. Specifically, for policy sequence $(\mu^{1},\ldots,\mu^{T})$ taken by an algorithm, the static regret is given, by
${\text{Reg}(T)}=\sum_{t=1}^{T}\langle\mu^{t},\ell_{t}\rangle-\min_{a\in\mathcal{A}}\sum^{T}_{t=1}\ell_{t}(a).$
$\mu^{t+1}(a)\propto\mu^{t}(a)e^{-\eta_{t}\ell_{t}(a)},\quad\text{for}\quad\forall a\in\mathcal{A}.$
where $\eta_{t}$ is the learning rate. See Algorithm 5 for the Hedge algorithm as applied to our problem setup.
A.2 $3$-player majority and minority game
In this section, we give a formal definition of the $3$-player majority and minority game.
We define the 3-player majority game as a symmetric zero-sum game with action space $\mathcal{A}:=\{0,1\}$ and the payoff function given by:
@@ A.2 $3$-player majority and minority game
Appendix B Proofs for Section 4
In the sequel, we will prove Proposition 4.1 in Section B.1, Proposition 4.2 in Section B.2 and Propostion 4.3 in Section B.3.
B.1 Proof of Proposition 4.1
In this section, we will prove (1), where both inequalities can be made strict in certain games.
@@ Appendix B Proofs for Section 4
To show the first inequality can be strict, we consider the $3$-player majority vote. Suppose $3$ players adopt the mixed strategies $(\alpha_{1},1-\alpha_{1})$, $(\alpha_{2},1-\alpha_{2})$ and $(\alpha_{3},1-\alpha_{3})$, respectively. It then holds that
$U_{1}(x_{1},x_{2},x_{3})$ $=U_{1}(\alpha_{1},\alpha_{2},\alpha_{3})$
$=\alpha_{1}\left(-(1-\alpha_{2})(1-\alpha_{3})+\frac{1}{2}\alpha_{2}(1-\alpha_{3})+\frac{1}{2}\alpha_{3}(1-\alpha_{2})\right)$
$\quad+(1-\alpha_{1})\left(-\alpha_{2}\alpha_{3}+\frac{1}{2}\alpha_{2}(1-\alpha_{3})+\frac{1}{2}\alpha_{3}(1-\alpha_{2})\right).$
$\max_{\alpha_{1}}\min_{\alpha_{2},\alpha_{3}}U_{1}(\alpha_{1},\alpha_{2},\alpha_{3})\leq\max_{\alpha_{1}}\min\{-\alpha_{1},-(1-\alpha_{1})\}=-\frac{1}{2}.$
Note that
$=\frac{1}{2}\min_{\alpha_{2},\alpha_{3}}\max\{-2(1-\alpha_{2})(1-\alpha_{3})+\alpha_{2}(1-\alpha_{3})+\alpha_{3}(1-\alpha_{2}),$
$\qquad\qquad\qquad\quad-2\alpha_{2}\alpha_{3}+\alpha_{2}(1-\alpha_{3})+\alpha_{3}(1-\alpha_{2})\}$
$=\frac{1}{2}\min_{\alpha_{2},\alpha_{3}}\max\{3(\alpha_{2}+\alpha_{3})-4\alpha_{2}\alpha_{3}-2,\alpha_{2}+\alpha_{3}-4\alpha_{2}\alpha_{3}\}$
$=0.$
Thus, we show that $\max_{\alpha_{1}}\min_{\alpha_{2},\alpha_{3}}U_{1}(\alpha_{1},\alpha_{2},\alpha_{3})<\min_{\alpha_{2},\alpha_{3}}\max_{\alpha_{1}}U_{1}(\alpha_{1},\alpha_{2},\alpha_{3})$, which implies the first inequality can be strict.
For the second inequality, due to a restriction on the minimization constraints, it is straightforward that
$\min_{x_{2},\cdots,x_{n}}\max_{x_{1}}U_{1}(x_{1},\cdots,x_{n})\leq\min_{x}\max_{x_{1}}U_{1}(x_{1},x^{\otimes n-1}).$
In the sequel, we prove $\min_{x}\max_{x_{1}}U_{1}(x_{1},x^{\otimes n-1})=0$ via contradiction. Note that by choosing $x_{1}=x$, we can show that
$\min_{x}\max_{x_{1}}U_{1}(x_{1},x^{\otimes n-1})\geq 0.$
@@
which contradicts with the assumption that the game is zero-sum and symmetric. Consequently, we prove the equation. As a result, we have
$\min_{x_{2},\cdots,x_{n}}\max_{x_{1}}U_{1}(x_{1},\cdots,x_{n})\leq\min_{x}\max_{x_{1}}U_{1}(x_{1},x^{\otimes n-1})=0.$
To show the second inequality can be strict, we consider a $3$-player minority game. If the other two players act 00 and $1$, respectively, then the learner always receive $-1/2$ payoff, which is strictly less than 00. We then finish the proofs. ∎
B.2 Proof of Proposition 4.2
In this section, we will prove (2), where the inequality can be made strict in certain games.
Proof of Proposition 4.2.
In the proof of Proposition 4.1, we have shown that $\min_{x}\max_{x_{1}}U_{1}(x_{1},x^{\otimes n-1})=0$. Thus, it remains to prove the inequality in (2).
@@
To show that the inequality can be strict, we consider the scenario where the learner is involved in a $3$-player majority game and plays a mixed strategy $(\beta,1-\beta)$ (i.e. play 00 w.p. $\beta$; play $1$ w.p. $1-\beta$). And the two opponents adopt an identical mixed strategy $(p,1-p)$ (i.e. play 00 w.p. $p$; play $1$ w.p. $1-p$). Then, we can calculate the payoff of the learner as $U_{1}(\beta,p,p)=\beta(-(1-p)^{2}+p(1-p))+(1-\beta)(-p^{2}+p(1-p))$. It then follows that $\max_{\beta\in[0,1]}\min_{p\in[0,1]}U_{1}(\beta,p,p)\leq\max_{\beta\in[0,1]}\min_{p\in\{0,1\}}U_{1}(\beta,p,p)=-1/2,$ which is strictly less than 00. Thus, we finish the proofs. ∎
B.3 Proof of Proposition 4.3
@@
Appendix C Proofs for Section 5
In Section C.1, we establish guarantees for the Hedge algorithm, $\mathrm{SAOL}^{\mathcal{H}}$, and behavior cloning. In Section C.2, we provide a detailed discussion of the matching lower bounds and prove Theorem 5.4 and Theorem 5.5.
C.1 Guarantees for efficient algorithms
In the sequel, we establish guarantees for the Hedge algorithm by proving Theorem 5.1 in Section C.1.1, for $\mathrm{SAOL}^{\mathcal{H}}$ by proving Theorem 5.2 in Section C.1.2, and for behavior cloning by proving Theorem 5.3 in Section C.1.3.
C.1.1 Proof of Theorem 5.1
In this section, we establish guarantees for the Hedge algorithm when facing fixed opponents.